# Efficient Algorithms for Finite $\mathbb{Z}$-Algebras

### v3

@@ Abstract.
For a finite $\mathbb{Z}$-algebra $R$, i.e., for a $\mathbb{Z}$-algebra which is a finitely generated $\mathbb{Z}$-module, we assume that $R$ is explicitly given by a system of $\mathbb{Z}$-module generators $G$, its relation module $\operatorname{Syz}(G)$, and the structure constants of the multiplication in $R$. In this setting we develop and analyze efficient algorithms for computing essential information about $R$. First we provide polynomial time algorithms for solving linear systems of equations over $R$ and for basic ideal-theoretic operations in $R$. Then we develop ZPP (zero-error probabilitic polynomial time) algorithms to compute the nilradical and the maximal ideals of 0-dimensional affine algebras $K[x_{1},\dots,x_{n}]/I$ with $K=\mathbb{Q}$ or $K=\mathbb{F}_{p}$. The task of finding the associated primes of a finite $\mathbb{Z}$-algebra $R$ is reduced to these cases and solved in ZPPIF (ZPP plus one integer factorization). With the same complexity, we calculate the connected components of the set of minimal associated primes $\operatorname{minAss}(R)$ and then the primitive idempotents of $R$. Finally, we prove that knowing an explicit representation of $R$ is polynomial time equivalent to knowing a strong Gröbner basis of an ideal $I$ such that $R=\mathbb{Z}[x_{1},\dots,x_{n}]/I$.
@@ 1. Introduction
Let us discuss the contents of the paper in more detail. Throughout we work with an explicitly given finite $\mathbb{Z}$-algebra $R$ defined as above. In Section 2 we start by using the well-known facts that the Smith and Hermite normal forms of an integer matrix can be calculated in polynomial time (see [19], [24], [31], [32]), in order to construct a polynomial time algorithm for solving linear systems of equations over $R$ (see Proposition 2.6). Based on this algorithm we perform various operations with ideals in $R$ efficiently, for instance ideal intersections (see Proposition 2.9). Moreover, we are able to compute preimages under the isomorphism given by the Chinese Remainder Theorem (see Proposition 2.10).
@@
In Section 3 we prepare later applications to $\mathbb{Z}$-algebras by reconsidering and reanalyzing some algorithms for 0-dimensional algebras over a field $K$. In order to compute the nilradical of an explicitly given $K$-algebra $R$, we need to calculate the factorization of univariate polynomials over $K$. The currently best algorithms have polynomial time complexity (P) in the case $K=\mathbb{Q}$ and zero-error probabilistic polynomial time complexity (ZPP) in the case of a prime field $\mathbb{F}_{p}$. Using such polynomial factorization algorithms, Algorithm 3.3 then determines the nilradical of $R$ with these time complexities (P resp. ZPP).
@@
Next we want to compute the primary decomposition of the zero ideal of $R$. In the case $K=\mathbb{F}_{p}$, we apply the method of Frobenius spaces (see [23], Alg. 5.2.7) and get an algorithm in ZPP (see Algorithm 3.6). In the case $K=\mathbb{Q}$, we can use the method of [23], Alg. 5.4.2 and get an algorithm in P. Altogether, by applying the primary decomposition algorithms to $R/\operatorname{Rad}(0)$, we are able to find the maximal ideals of $R$ in P and ZPP for $K=\mathbb{Q}$ and $K=\mathbb{F}_{p}$, respectively (see Corollary 3.8).
@@
In Section 4 we then use the ideas of [27] to compute the associated primes of an explicitly given finite $\mathbb{Z}$-algebra $R$ (see Algorithm 4.2). The method is to distinguish between the prime ideals which contain an integer prime number and those which don’t. In both cases the computation is reduced to the setting of the preceding section, i.e., to computations of 0-dimensional algebras over fields. Since the determination of associated primes may involve the factorization of an integer, the best time complexity we can achieve here is ZPPIF, i.e., ZPP plus one integer factorization. More precisely, the integer which has to be factored is the torsion exponent of the additive group of $R$.
@@
In Section 5 we treat the next topics, namely the computation of the primitive idempotents of an explicitly given finite $\mathbb{Z}$-algebra $R$ and the connected components of its prime spectrum. It turns out that it is advisable to solve the latter task first. More precisely, since $R$ may have infinitely many prime ideals, we compute the connected components of the set of minimal associated primes $\operatorname{minAss}(R)$ in Algorithm 5.6. This algorithm uses the results of the preceding section, whence its worst-case time complexity is in ZPPIF. Finally, we calculate the primitive idempotents of $R$ in Algorithm 5.8 by lifting them from the primitive idempontents of $R/\operatorname{Rad}(0)$. This lifting process is performed using Algorithm 5.1. Once again the worst-case time complexity is in ZPPIF.
@@
In the last section we connect the method of using an explicit representation of $R$ to the more traditional method of calculating a strong Gröbner basis, when $R$ is given as $R=P/I$ with $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ and an ideal $I$ in $P$ whose generators are known. Starting with an explicitly given finite $\mathbb{Z}$-algebra $R$, we can compute in polynomial time a strong Gröbner basis of a defining ideal $I$ of $R$ (see Corollary 6.5). For this direction we use a generalization of the Buchberger-Möller Algorithm (see Algorithm 6.3). Conversely, starting with a strong Gröbner basis of $I$, we can calculate an explicit representation of $R$ in polynomial time (see Algorithm 6.7). For this direction, we use a generalization of Macaulay’s Basis Theorem to finite $\mathbb{Z}$-algebras (see Proposition 6.6).
@@ 2. Polynomial Time Computations in Finite $\mathbb{Z}$-Algebras
Proof.
The product of the invariant factors of $R^{+}$ is given by the gcd of all maximal rank minors of $A$. The torsion exponent is therefore bounded by the absolute value of a non-zero maximal rank minor of $A$. Hadamard’s inequality then yields $\tau\leq n^{n/2}\left\lVert A\right\rVert$, which means the bit complexity of $\tau$ is bounded by $n\log_{2}(n\left\lVert A\right\rVert)$. ∎
@@
(ii) $\sum\limits_{k=1}^{p}\textstyle\sum\limits_{i,j,\ell=0}^{n}z_{ki}b_{kj}c_{ij\ell}e_{\ell}-\textstyle\sum\limits_{j=1}^{m}\textstyle\sum\limits_{i=0}^{n}w_{j}a_{ij}e_{i}\;=\;0$
@@
This is the case if and only if there exist $\alpha_{1},\dots,\alpha_{m}\in\mathbb{Z}$ such that the left hand side is equal to $\sum_{j=1}^{m}\sum_{i=1}^{n}\alpha_{j}a_{ij}g_{i}$. The claim then follows by rewriting the products $g_{i}g_{j}$ using the structure constants of $R$ and applying the canonical isomorphism $R\cong\mathbb{Z}^{n+1}/\operatorname{Syz}(G)$. ∎
@@
over $R$. Every solution is of the form $(h_{1},h_{2},h_{3})\in R^{3}$ with $h_{i}=d_{i1}g_{1}+d_{i2}g_{2}+d_{i3}g_{3}$, where $d_{ij}\in\mathbb{Z}$. To compute generators of the solution space $\mathcal{L}\subseteq R^{3}$, we follow Proposition 2.6 and substitute $x_{i}$ by $h_{i}$. Using the structure constants, we then replace products $g_{i}g_{j}$ by $\mathbb{Z}$-linear combinations of the generators and obtain the system of equations
@@
Proposition 2.6 yields the following complexity bound for solving linear equations.
Corollary 2.8.
Let $R$ be an explicitly given finite $\mathbb{Z}$-algebra. Generators of the solution space of a linear equation over $R$ as in Proposition 2.6 can be computed in polynomial time in the bit complexity of the input which is given by $\gamma$ (for $R$) and by $p(n+1)\log_{2}(M)$ where $M=\max\{b_{kj}\}$ (for the elements $f_{1},\dots,f_{p})$.
Proof.
The coefficients in the system of equations (ii) in Proposition 2.6 are $b_{kj}$, $c_{ij\ell}$ and $a_{ij}$. The claim therefore follows immediately from Remark 2.4.b. ∎
@@
Proof.
To prove (a), let $A=(a_{ij})\in\operatorname{Mat}_{n+1,m}(\mathbb{Z})$ be the matrix whose rows are given by the generators of $\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})$. The rank and the invariant factors of $R^{+}$ can be determined from the Smith normal form of $A$, which can be computed in polynomial time by Remark 2.4.
@@
Generators of $\operatorname{Syz}_{R}(M)$ can be computed in polynomial time by solving an appropriate linear system of equations over $R$ using Proposition 2.6. The first non-zero coordinates of the generators then generate $J\cap J^{\prime}$ by [21], Proposition 3.2.3. ∎
@@
    Using Proposition 2.9.b, compute $\mathbb{Z}$-submodules $V_{j}\subseteq\mathbb{Z}^{n+1}$ such that we have $R/J_{j}\;\cong\;\mathbb{Z}^{n+1}/V_{j}$ for $j=1,\dots,s$.
@@
Proof.
The tuple $h$ satisfies $h\in\bigcap_{j\neq i}V_{j}$ and $h-(1,0,\dots,0)\in V_{i}$. This shows that the residue class of $h$ in $\mathbb{Z}^{n+1}/\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})$ is mapped to $e_{i}$ under the canonical map $\psi:\;\mathbb{Z}^{n+1}/\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})\longrightarrow\mathbb{Z}^{n+1}/V_{1}\times\cdots\times\mathbb{Z}^{n+1}/V_{s}$. Hence $f$ is mapped to $e_{i}$ under the map $\varphi$. Steps (1) and (2) of the algorithm can be performed in polynomial time by Proposition 2.9. The linear system in Step (3) can also be solved in polynomial time by Remark 2.4. ∎
@@
    Using Proposition 2.9.b, we find $V_{1}=\langle(1,0,1,0),(0,0,0,2),(-1,1,0,0)\rangle$ and $V_{2}=\langle(0,0,3,0),(0,1,0,0),(0,0,0,1)\rangle$ such that $R/J_{1}\cong\mathbb{Z}^{4}/V_{1}$ and $R/J_{2}\cong\mathbb{Z}^{4}/V_{2}$.
@@ 3. Computing the Maximal Ideals of a 0-Dimensional $K$-Algebra
The bit complexity of the presentation of $R/\operatorname{Rad}(0)$ is polynomially bounded by the bit complexity of the input, since during each iteration the bit complexity of the structure constants obtained in step (6) is polynomially bounded by the bit complexity of the structure constants of the previous iteration. ∎
@@
Proof.
The correctness of this algorithm is shown in [23], Algorithm 5.2.11. In particular, it is proved there that $t=d_{1}+\cdots+d_{u}$ throughout the course of this algorithm. Therefore the number of iterations of steps (3)–(6) is bounded by $s$ which in turn is bounded by the vector space dimension $n$ of $R$. Algorithm 5.2.7 in step (4) involves computing a basis for the kernel of a matrix over $K$ and can therefore be done in polynomial time. As discussed in the proof of Algorithm 3.3, the minimal polynomial in step (5) can also be computed in polynomial time. Computing its factorization is in ZPP by Remark 3.2. ∎
@@
Proof.
Using Algorithm 3.3, we compute the nilradical $\operatorname{Rad}(0)$ of $R$ in polynomial time. This algorithm also yields an explicit presentation of $R/\operatorname{Rad}(0)$. If $K=\mathbb{Q}$, we then apply Algorithm 7.2 from [25] to $R/\operatorname{Rad}(0)$ and obtain the maximal ideals of $R$ in polynomial time. Similarly, in the case $K=\mathbb{F}_{p}$, we apply Algorithm 3.6 to $R/\operatorname{Rad}(0)$. ∎
@@ 4. Computing the Associated Primes of Finite $\mathbb{Z}$-Algebras
Proof.
Part (a) follows immediately from the definition of the exponent of the torsion subgroup of $R^{+}$. It then implies $I:\langle\tau\rangle=I:\langle\tau\rangle^{\infty}$, which means that claim (b) is a standard lemma in commutative algebra. To prove (c), we note that the ring $P/I$ is finite if and only if there exists a positive integer $k\in\mathbb{Z}$ with $I\cap\mathbb{Z}=\langle k\rangle$. If such a number $k$ exists, we have $k\cdot f=0$ for all $f\in R$, and therefore $\tau\mid k$. But we also have $n\cdot 1=0$ in $R=P/I$, and hence $\tau\in I$. This implies $k\mid\tau$, and thus $k=\tau$. ∎
@@
Proof.
The correctness of the algorithm follows from Lemma 4.1 and Proposition 4.7 in [20]. Let us analyze the complexity of each of its steps. The torsion exponent and the rank of $R$ can be computed in polynomial time in $\beta$ using Proposition 2.9.a, and the bit complexity of the torsion exponent is polynomially bounded by Lemma 2.5. Since $P/I$ is a finite $\mathbb{Z}$-algebra, the ideals $I\mathbb{Q}[x_{1},\dots,x_{n}]$ and $I\mathbb{F}_{p}[x_{1},\dots,x_{n}]$ are 0-dimensional and therefore define 0-dimensional $\mathbb{Q}$- and $\mathbb{F}_{p}$-algebras, respectively. Their vector space dimension is less than or equal to the number of generators of $R$, and their structure constants are given by the structure constants of $R$. Thus we obtain the maximal components in lines (4) and (14) in polynomial and probabilistic polynomial time, respectively, by applying the algorithms in Section 3. The intersection of the prime ideals in line (7) can be computed in polynomial time by Proposition 2.9.e. Finally, Proposition 2.9.c allows us to check the containment of ideals in line (8) in polynomial time.
In summary, all steps except for the integer prime factorization in line (11) are in ZPP. ∎
@@
    Using Proposition 2.9.a, we find that the torsion exponent of $R$ is 6.
@@
    Here we determine the minimal associated primes $\bar{\mathfrak{p}}_{3}=\langle x+1,y,z\rangle$ and $\bar{\mathfrak{p}}_{4}=\langle x,y,z\rangle$ of $I\mathbb{F}_{3}[x,y,z]$ using Algorithm 3.6.
@@
    We compute $J=\mathfrak{p}_{1}\cap\mathfrak{p}_{2}=\langle z,y,3x-3\rangle$.
@@
    The ideal $J$ is contained in both $\mathfrak{p}_{3}$ and $\mathfrak{p}_{4}$.

### v5

@@ Abstract.
For a finite $\mathbb{Z}$-algebra $R$, i.e., for a $\mathbb{Z}$-algebra which is a finitely generated $\mathbb{Z}$-module, we assume that $R$ is explicitly given by a system of $\mathbb{Z}$-module generators $G$, its relation module $\operatorname{Syz}(G)$, and the structure constants of the multiplication in $R$. In this setting we develop and analyze efficient algorithms for computing essential information about $R$. First we provide polynomial time algorithms for solving linear systems of equations over $R$ and for basic ideal-theoretic operations in $R$. Then we develop ZPP (zero-error probabilitic polynomial time) algorithms to compute the nilradical and the maximal ideals of 0-dimensional affine algebras $K[x_{1},\dots,x_{n}]/I$ with $K=\mathbb{Q}$ or $K=\mathbb{F}_{p}$. The task of finding the associated primes of a finite $\mathbb{Z}$-algebra $R$ is reduced to these cases and solved in ZPPIF (ZPP plus one integer factorization). With the same complexity, we calculate the connected components of the set of minimal associated primes $\operatorname{minPrimes}(R)$ and then the primitive idempotents of $R$. Finally, we prove that knowing an explicit representation of $R$ is polynomial time equivalent to knowing a strong Gröbner basis of an ideal $I$ such that $R=\mathbb{Z}[x_{1},\dots,x_{n}]/I$.
@@ 1. Introduction
Let us discuss the contents of the paper in more detail. Throughout we work with an explicitly given finite $\mathbb{Z}$-algebra $R$ defined as above. In Section 2 we start by using the well-known facts that the Smith and Hermite normal forms of an integer matrix can be calculated in polynomial time (see [19], [24], [31], [32]), in order to construct a polynomial time algorithm for solving linear systems of equations over $R$ (see Proposition 5). Based on this algorithm we perform various operations with ideals in $R$ efficiently, for instance ideal intersections (see Proposition 7). Moreover, we are able to compute preimages under the isomorphism given by the Chinese Remainder Theorem (see Proposition 8).
@@
In Section 3 we prepare later applications to $\mathbb{Z}$-algebras by reconsidering and reanalyzing some algorithms for 0-dimensional algebras over a field $K$. In order to compute the nilradical of an explicitly given $K$-algebra $R$, we need to calculate the factorization of univariate polynomials over $K$. The currently best algorithms have polynomial time complexity (P) in the case $K=\mathbb{Q}$ and zero-error probabilistic polynomial time complexity (ZPP) in the case of a prime field $\mathbb{F}_{p}$. Using such polynomial factorization algorithms, Algorithm 10 then determines the nilradical of $R$ with these time complexities (P resp. ZPP).
@@
Next we want to compute the primary decomposition of the zero ideal of $R$. In the case $K=\mathbb{F}_{p}$, we apply the method of Frobenius spaces (see [23], Alg. 5.2.7) and get an algorithm in ZPP (see Algorithm 11). In the case $K=\mathbb{Q}$, we can use the method of [23], Alg. 5.4.2 and get an algorithm in P. Altogether, by applying the primary decomposition algorithms to $R/\operatorname{Rad}(0)$, we are able to find the maximal ideals of $R$ in P and ZPP for $K=\mathbb{Q}$ and $K=\mathbb{F}_{p}$, respectively (see Corollary 12).
@@
In Section 4 we then use the ideas of [27] to compute the associated primes of an explicitly given finite $\mathbb{Z}$-algebra $R$ (see Algorithm 14). The method is to distinguish between the prime ideals which contain an integer prime number and those which don’t. In both cases the computation is reduced to the setting of the preceding section, i.e., to computations of 0-dimensional algebras over fields. Since the determination of associated primes may involve the factorization of an integer, the best time complexity we can achieve here is ZPPIF, i.e., ZPP plus one integer factorization. More precisely, the integer which has to be factored is the torsion exponent of the additive group of $R$.
@@
In Section 5 we treat the next topics, namely the computation of the primitive idempotents of an explicitly given finite $\mathbb{Z}$-algebra $R$ and the connected components of its prime spectrum. It turns out that it is advisable to solve the latter task first. More precisely, since $R$ may have infinitely many prime ideals, we compute the connected components of the set of minimal associated primes $\operatorname{minPrimes}(R)$ in Algorithm 18. This algorithm uses the results of the preceding section, whence its worst-case time complexity is in ZPPIF. Finally, we calculate the primitive idempotents of $R$ in Algorithm 19 by lifting them from the primitive idempontents of $R/\operatorname{Rad}(0)$. This lifting process is performed using Algorithm 15. Once again the worst-case time complexity is in ZPPIF.
@@
In the last section we connect the method of using an explicit representation of $R$ to the more traditional method of calculating a strong Gröbner basis, when $R$ is given as $R=P/I$ with $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ and an ideal $I$ in $P$ whose generators are known. Starting with an explicitly given finite $\mathbb{Z}$-algebra $R$, we can compute in polynomial time a strong Gröbner basis of a defining ideal $I$ of $R$ (see Corollary 22). For this direction we use a generalization of the Buchberger-Möller Algorithm (see Algorithm 21). Conversely, starting with a strong Gröbner basis of $I$, we can calculate an explicit representation of $R$ in polynomial time (see Algorithm 24). For this direction, we use a generalization of Macaulay’s Basis Theorem to finite $\mathbb{Z}$-algebras (see Proposition 23).
@@ 2. Polynomial Time Computations in Finite $\mathbb{Z}$-Algebras
Proof 2.1.
The product of the invariant factors of $R^{+}$ is given by the gcd of all maximal rank minors of $A$. The torsion exponent is therefore bounded by the absolute value of a non-zero maximal rank minor of $A$. Hadamard’s inequality then yields $\tau\leq n^{n/2}\left\lVert A\right\rVert$, which means the bit complexity of $\tau$ is bounded by $n\log_{2}(n\left\lVert A\right\rVert)$.
@@
$\sum\limits_{k=1}^{p}\textstyle\sum\limits_{i,j,\ell=0}^{n}z_{ki}b_{kj}c_{ij\ell}e_{\ell}-\textstyle\sum\limits_{j=1}^{m}\textstyle\sum\limits_{i=0}^{n}w_{j}a_{ij}e_{i}\;=\;0$ (ii)
@@
This is the case if and only if there exist $\alpha_{1},\dots,\alpha_{m}\in\mathbb{Z}$ such that the left hand side is equal to $\sum_{j=1}^{m}\sum_{i=1}^{n}\alpha_{j}a_{ij}g_{i}$. The claim then follows by rewriting the products $g_{i}g_{j}$ using the structure constants of $R$ and applying the canonical isomorphism $R\cong\mathbb{Z}^{n+1}/\operatorname{Syz}(G)$.
@@
over $R$. Every solution is of the form $(h_{1},h_{2},h_{3})\in R^{3}$ with $h_{i}=d_{i1}g_{1}+d_{i2}g_{2}+d_{i3}g_{3}$, where $d_{ij}\in\mathbb{Z}$. To compute generators of the solution space $\mathcal{L}\subseteq R^{3}$, we follow Proposition 5 and substitute $x_{i}$ by $h_{i}$. Using the structure constants, we then replace products $g_{i}g_{j}$ by $\mathbb{Z}$-linear combinations of the generators and obtain the system of equations
@@
Proposition 5 yields the following complexity bound for solving linear equations.
Corollary 6.
Let $R$ be an explicitly given finite $\mathbb{Z}$-algebra. Generators of the solution space of a linear equation over $R$ as in Proposition 5 can be computed in polynomial time in the bit complexity of the input which is given by $\gamma$ (for $R$) and by $p(n+1)\log_{2}(M)$ where $M=\max\{b_{kj}\}$ (for the elements $f_{1},\dots,f_{p})$.
Proof 2.3.
The coefficients in the system of equations (ii) in Proposition 5 are $b_{kj}$, $c_{ij\ell}$ and $a_{ij}$. The claim therefore follows immediately from Remark 3.b.
@@
Proof 2.4.
To prove (a), let $A=(a_{ij})\in\operatorname{Mat}_{n+1,m}(\mathbb{Z})$ be the matrix whose rows are given by the generators of $\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})$. The rank and the invariant factors of $R^{+}$ can be determined from the Smith normal form of $A$, which can be computed in polynomial time by Remark 3.
@@
Generators of $\operatorname{Syz}_{R}(M)$ can be computed in polynomial time by solving an appropriate linear system of equations over $R$ using Proposition 5. The first non-zero coordinates of the generators then generate $J\cap J^{\prime}$ by [21], Proposition 3.2.3.
@@
    Using Proposition 7.b, compute $\mathbb{Z}$-submodules $V_{j}\subseteq\mathbb{Z}^{n+1}$ such that we have $R/J_{j}\;\cong\;\mathbb{Z}^{n+1}/V_{j}$ for $j=1,\dots,s$.
@@
Proof 2.5.
The tuple $h$ satisfies $h\in\bigcap_{j\neq i}V_{j}$ and $h-(1,0,\dots,0)\in V_{i}$. This shows that the residue class of $h$ in $\mathbb{Z}^{n+1}/\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})$ is mapped to $e_{i}$ under the canonical map $\psi:\;\mathbb{Z}^{n+1}/\operatorname{Syz}_{\mathbb{Z}}(\mathcal{G})\longrightarrow\mathbb{Z}^{n+1}/V_{1}\times\cdots\times\mathbb{Z}^{n+1}/V_{s}$. Hence $f$ is mapped to $e_{i}$ under the map $\varphi$. Steps (1) and (2) of the algorithm can be performed in polynomial time by Proposition 7. The linear system in Step (3) can also be solved in polynomial time by Remark 3.
@@
    Using Proposition 7.b, we find $V_{1}=\langle(1,0,1,0),(0,0,0,2),(-1,1,0,0)\rangle$ and $V_{2}=\langle(0,0,3,0),(0,1,0,0),(0,0,0,1)\rangle$ such that $R/J_{1}\cong\mathbb{Z}^{4}/V_{1}$ and $R/J_{2}\cong\mathbb{Z}^{4}/V_{2}$.
@@ 3. Computing the Maximal Ideals of a 0-Dimensional $K$-Algebra
The bit complexity of the presentation of $R/\operatorname{Rad}(0)$ is polynomially bounded by the bit complexity of the input, since during each iteration the bit complexity of the structure constants obtained in step (6) is polynomially bounded by the bit complexity of the structure constants of the previous iteration.
@@
Proof 3.2.
The correctness of this algorithm is shown in [23], Algorithm 5.2.11. In particular, it is proved there that $t=d_{1}+\cdots+d_{u}$ throughout the course of this algorithm. Therefore the number of iterations of steps (3)–(6) is bounded by $s$ which in turn is bounded by the vector space dimension $n$ of $R$. Algorithm 5.2.7 in step (4) involves computing a basis for the kernel of a matrix over $K$ and can therefore be done in polynomial time. As discussed in the proof of Algorithm 10, the minimal polynomial in step (5) can also be computed in polynomial time. Computing its factorization is in ZPP by Remark 9.
@@
Proof 3.3.
Using Algorithm 10, we compute the nilradical $\operatorname{Rad}(0)$ of $R$ in polynomial time. This algorithm also yields an explicit presentation of $R/\operatorname{Rad}(0)$. If $K=\mathbb{Q}$, we then apply Algorithm 7.2 from [25] to $R/\operatorname{Rad}(0)$ and obtain the maximal ideals of $R$ in polynomial time. Similarly, in the case $K=\mathbb{F}_{p}$, we apply Algorithm 11 to $R/\operatorname{Rad}(0)$.
@@ 4. Computing the Associated Primes of Finite $\mathbb{Z}$-Algebras
Proof 4.1.
Part (a) follows immediately from the definition of the exponent of the torsion subgroup of $R^{+}$. It then implies $I:\langle\tau\rangle=I:\langle\tau\rangle^{\infty}$, which means that claim (b) is a standard lemma in commutative algebra. To prove (c), we note that the ring $P/I$ is finite if and only if there exists a positive integer $k\in\mathbb{Z}$ with $I\cap\mathbb{Z}=\langle k\rangle$. If such a number $k$ exists, we have $k\cdot f=0$ for all $f\in R$, and therefore $\tau\mid k$. But we also have $n\cdot 1=0$ in $R=P/I$, and hence $\tau\in I$. This implies $k\mid\tau$, and thus $k=\tau$.
@@
Proof 4.2.
The correctness of the algorithm follows from Lemma 13 and Proposition 4.7 in [20]. Let us analyze the complexity of each of its steps. The torsion exponent and the rank of $R$ can be computed in polynomial time in $\beta$ using Proposition 7.a, and the bit complexity of the torsion exponent is polynomially bounded by Lemma 4. Since $P/I$ is a finite $\mathbb{Z}$-algebra, the ideals $I\mathbb{Q}[x_{1},\dots,x_{n}]$ and $I\mathbb{F}_{p}[x_{1},\dots,x_{n}]$ are 0-dimensional and therefore define 0-dimensional $\mathbb{Q}$- and $\mathbb{F}_{p}$-algebras, respectively. Their vector space dimension is less than or equal to the number of generators of $R$, and their structure constants are given by the structure constants of $R$. Thus we obtain the maximal components in lines (4) and (14) in polynomial and probabilistic polynomial time, respectively, by applying the algorithms in Section 3. The intersection of the prime ideals in line (7) can be computed in polynomial time by Proposition 7.e. Finally, Proposition 7.c allows us to check the containment of ideals in line (8) in polynomial time.
In summary, all steps except for the integer prime factorization in line (11) are in ZPP.
@@
    Using Proposition 7.a, we find that the torsion exponent of $R$ is 6.
@@
    Here we calculate the prime factorization $6=2\cdot 3$.
  6. (14)
    We determine the minimal associated primes $\bar{\mathfrak{p}}_{3}=\langle x+1,y,z\rangle$ and $\bar{\mathfrak{p}}_{4}=\langle x,y,z\rangle$ of $I\mathbb{F}_{3}[x,y,z]$ using Algorithm 11.
@@
    We determine the minimal associated primes $\bar{\mathfrak{p}}_{5}=\langle x+1,y,z\rangle$ and $\bar{\mathfrak{p}}_{6}=\langle x,y,z\rangle$ of $I\mathbb{F}_{2}[x,y,z]$ using Algorithm 11.
@@
    Their canonical liftings are $\mathfrak{p}_{5}=\langle x+1,y,z,2\rangle$ and $\mathfrak{p}_{6}=\langle x,y,z,2\rangle$.
  10. (7)
    We compute $J=\mathfrak{p}_{1}\cap\mathfrak{p}_{2}=\langle z,y,x^{2}+x-6\rangle$.
  11. (8)
    The ideal $J$ is contained in $\mathfrak{p}_{3}$, $\mathfrak{p}_{4}$, $\mathfrak{p}_{5}$, and $\mathfrak{p}_{6}$.